Consider the following information:

  

Rate of Return if State Occurs

  State of

Probability of

  Economy

State of Economy

Stock A

Stock B

Stock C

  Boom

0.58

0.07

0.15

0.33

  Bust

0.42

0.16

0.06

0.06

  

a.

What is the expected return on an equally weighted portfolio of these three stocks? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16))

  Expected return

%

  

b.

What is the variance of a portfolio invested 20 percent each in A and B and 60 percent in C? (Do not round intermediate calculations and round your answer to 6 decimal places. (e.g., 32.161616))

  Variance
Answer #1

page- calculation of expected refuen prob stolt A stock 15 stolle c 0.58 0.42 0.04 0.16 0.15 0.06 0.33 -0.06 expected returnpage- portfolio variance formula = WA TATWB Bt Wetet quA WB. CONA,B + 2HB. We coupe + 2 We. WA CONCA TA = Je prob.da dA Boomprob ore Å dare 0.C Boom Bust & 0.7 0.17 033 -0.06 0.42 08529 -0.23 se = 0.56*0.0206+042 30-0529 Jo.037066 Tre = 0.192505 CONWA = 20) WB = 30). We = 60). Variance = WA TAA WR. Btwee of 2. Wp -Cola, * 2 Woche comes of g We Wa colea = (0,20x 0.044474)